Find the downward force applied 25 cm from the axis of rotation when the torque is found to be 23.75 Nm


1
Expert's answer
2021-01-16T17:22:47-0500

By definition, the torque "\tau" is the product of the force "F" and the distance of the force application "r" (if they are perpendicular). Thus, oobtain:


"\tau = Fr"

Since "\tau = 23.75\space N\cdot m", and "r = 0.25m", obtain the force:


"F = \dfrac{\tau}{r} = \dfrac{23.75\space N\cdot m}{0.25m} = 95N"

Answer. 95 N.